OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4 My answer is y = x - 4 Am I right? :)

OpenStudy (anonymous):

yes that should be correct good job

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

I like your username too, lol :)

OpenStudy (anonymous):

i take that back it's not here's the formula y =mx+b is slope-intercept form...m is slope

OpenStudy (anonymous):

and the rule is that if lines r parallel they have the same slope....so

OpenStudy (anonymous):

Is it y = -x - 2 then?

OpenStudy (anonymous):

so u where right sorry for confusing u i just was checking

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

aw man i messed up again it's the same slope as the first equation

OpenStudy (anonymous):

Lol... So which one is it? y = x - 4 or y = -x - 2

OpenStudy (anonymous):

so witch is the last unit is b=y-intercept.....so what is the y-intercept....then that will be the answer....sorry i can't tell u.....but i can confirm after you pick one

OpenStudy (anonymous):

y = x - 4?

OpenStudy (anonymous):

That makes more sense, thank you so much for taking your time to help me! :)

OpenStudy (anonymous):

aw man i'm so sorry i did it again it's y = -x - 2

OpenStudy (anonymous):

lol... Are you messing with me?

OpenStudy (anonymous):

it's that cause the slope stays the same and if u construct it correctly it will decline and when it declines it passes through the negative 2

OpenStudy (anonymous):

so y=mx+b so m=slope=the same as the original=-1or in this case=-x and b=y-intercept= which means it's where the line passes through the y-axis so b=-2......so sorry for being confusing

OpenStudy (anonymous):

@hunterhoff

OpenStudy (anonymous):

So the answer is y = -x - 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Its fine! :) Im no more confused than you... lol

OpenStudy (anonymous):

Thanks angain! :)

OpenStudy (anonymous):

np